# Partial Identification Using Neural Causal Models with Continuous Treatment

### v2

@@ 3 Approximation Error of Neural Causal Models
In this section, we study the expressive power of NCMs, which serves as a key ingredient in proving the consistency result. In particular, given an SCM $\mathcal{M}^{*}$, we want to construct an NCM $\widehat{\mathcal{M}}$ such that the two causal models produce similar interventional results. Unlike in the discrete case [4, 51], latent distributions can be extremely complicated in general cases. The main challenge is how to design the structure of NCMs to ensure strong approximation power.
@@ A SCM $\mathcal{M}$ with causal graph $\mathcal{G}$ has canonical form if
$\widehat{V}_{i}=\begin{cases}\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right),&V_{i}\text{ is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}+\log\left(\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right)\right)_{k}\right\},&V_{i}\text{ is categorical},\end{cases}$ (3.3)
wheres $N_{C,j}$ are constants to be specified later.
For the first part, we need to approximate Lipschitz continuous functions. For simplicity, we assume that the domain of the functions are uniform cubes. Similar arguments hold for any bounded cubes. We denote $\mathcal{NN}_{k_{1},k_{2}}(W,L)$ to be the set of ReLu NNs with input dimension $k_{1}$, output dimension $k_{2}$, width $W$ and depth $L$. It has been shown that $\epsilon(\mathcal{NN}_{k_{1},1}(2d_{1}+10,L_{0}),\mathcal{F}_{L}([0,1]^{d_{1}},\mathbb{R}))\leqslant O(L_{0}^{-{2}/{k_{1}}})$ [53], where $\epsilon(\cdot,\cdot)$ denotes the approximation error defined in Section 2. For a vector valued function, we can use a wider NN to approximate each coordinate and get a similar rate.
@@
3.1 Approximating Mixed Distributions by Wide Neural Networks
@@
Figure 1: Architecture of wide neural network for $4-$dimensional output. The first (yellow) part approximates the distribution on different connected components of the support using the results from [43]. The width and depth of each block in this part are $W_{1}$ and $L_{1}$. The second (blue) part transforms the distributions on the unit cube to the distributions on the support. The width and depth of each block in the blue part are $W_{2}$ and $L_{2}$. The third (green) part is the Gumbel-Softmax layer. It combines the distributions on different connected components of the support together and outputs the final distribution.
@@
In this subsection, we will extend the results in [43] to construct a wide neural network as the push-forward map. It turns out that to get a good approximation of the targeted distribution, the shape of the support is essential.
@@
3.4 encompasses almost all natural distributions. For example, distributions supported on $[0,1]^{d}$ and closed balls, finitely supported distributions and mixtures of them all satisfy this assumption. 3.4 allows us to transform the support of the targeted distribution into unit cubes and the nice geometric properties of unit cubes facilitate our construction.
Now, we briefly explain the construction of the push-forward maps. An example is provided in Section B.1 to illustrate the construction. The NN architecture consists of three parts (see Figure 2). The input dimension is the same as the number of connected components of the support $N_{C}$. For each component $C_{i}$, let $H_{i}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$, where $d_{i}^{C}$ is the dimension of component $C_{i}$ in 3.4. By $\mathbb{P}=(H_{i})_{\#}(H_{i}^{-1})_{\#}\mathbb{P}$ on $C_{i}$, we can approximate $(H_{i}^{-1})_{\#}\mathbb{P}$ first, which is supported on a unit cube. [43] constructs a wide NN $\widehat{g}$ of width $W$ and constant depth such that $W(\widehat{g}_{\#}\lambda,(H_{i})^{-1}_{\#}\mathbb{P})\leqslant O(W^{-1/d_{i}^{C}})$ where $\lambda$ is the uniform measure on $[0,1]$. Then, we approximate the Lipschitz map $H_{i}$ to $C_{i}$ to pull the distribution back to $C_{i}$. These are the first two parts (yellow and blue blocks in Figure 2) of the architecture.
Suppose that the output of $i$-th coordinate in the first two parts is $\vec{v}_{i}$, the Gumbel-softmax layer in the third part (green box) combines different components of the support. In particular, we want to output $\vec{v}_{i}$ with probability $p_{i}=\mathbb{P}(C_{i})$. Let $V=[\vec{v}_{1},\cdots,\vec{v}_{i}]$, this can be achieved by outputting $VX$, where $X=(X_{1},\cdots,X_{N_{C}})^{\mathsf{T}}$ is a one-hot random vector with $\mathbb{P}({X}_{i}=1)=p_{i}$. To use backpropagation in training, we use the Gumbel-Softmax trick [31] to (approximately) simulate such a random vector, $\widehat{X}^{\tau}_{i}=\frac{\exp((\log p_{i}+G_{i})/\tau)}{\sum_{k=1}^{N_{C}}\exp((\log p_{k}+G_{k})/\tau)},$ where $\tau>0$ is the temperature parameter (a hyperparameter) and $G_{i}\sim\text{Gumbel}(0,1)$ are i.i.d. standard Gumbel variables. As $\tau\rightarrow 0$, the distribution of $\widehat{X}^{\tau}$ converges almost surely to the categorical distribution [36, Proposition 1]. In particular, when $\tau=0$, we denote $\widehat{X}_{i}^{0}=\mathbb{I}_{i=\arg\max_{j}\{\log p_{i}+G_{i}\}}$. The output of the last layer is $V\widehat{X}^{\tau}$. Note that the Gumbel-softmax function is differentiable with respect to parameter $\{\log(p_{i})\}_{i=1,\cdots,N_{C}}$. Therefore, we can train the network with common gradient-based algorithms. Putting things together, we obtain the following theorem.
Theorem 3.5.
Given any probability measure $\mathbb{P}$ on $\mathbb{R}^{d}$ that satisfies 3.4, let $\lambda$ be the Lebesgue measure on $[0,1]^{N_{C}}$, where $N_{C}$ is defined in 3.4. There exists a neural network $\widehat{g}=\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}$ with the above architectsure (Figure 2) such that
Here, $\widehat{g}_{i},i=1,2$ has the separable form $\left(\widehat{g}_{i}^{1}(x_{1}),\cdots,\widehat{g}_{i}^{N_{C}}(x_{N_{C}})\right)$ and $\widehat{g}_{1}^{j}\in\mathcal{NN}_{1,d_{j}^{C}}(W_{1},\Theta(d_{j}^{C}))$, $\ \widehat{g}_{2}^{j}\in\mathcal{NN}_{d_{j}^{C},d}(\Theta(d\cdot d_{j}^{C}),L_{2}),\ j\in[N_{C}]$. $\{d_{j}^{C}\}$ are the dimension of cubes in 3.4. $\widehat{g}_{3}^{\tau}$ is the Gumbel-Softmax layer with temperature parameter $\tau>0$.
3.2 Approximating Mixed Distributions by Deep Neural Networks
In this subsection, we will show that under one additional assumption on the distribution, deep ReLu networks have a stronger approximation power in approximating distributions, which means we can use fewer computational units to achieve the same worst-case theoretical approximation error.
@@ 3.2 Approximating Mixed Distributions by Deep Neural Networks
3.6 implies that $d\lambda/d(f^{-1}_{\#}\mathbb{P})$ exists and is lowered bounded by a constant $C_{f}$. The next proposition extends the Skorohod representation theorem [48]. It shows that under 3.6, it is possible to simulate any distribution on the unit cubes with Hölder continuous curves and uniform distribution on $[0,1]$.
Proposition 3.7.
Let $\lambda$ be the Lebesgue measure on $[0,1]$. Given any probability measure $\mathbb{P}$ that satisfies 3.6, there exists a continuous curve $\gamma:[0,1]\rightarrow\text{supp}(\mathbb{P})$ such that $\gamma_{\#}\lambda=\mathbb{P}$. Furthermore, if $d\geqslant 1$, $\gamma$ is $1/d$-Hölder continuous.
Results from [53] show that we can approximate any Hölder continuous $d$-dimensional function with index $\alpha$ by a deep ReLu network with depth $L$ and error $O(L^{-2\alpha/d})$. Leveraging this result, we can replace the first part of the architecture in the previous subsection with deep ReLu networks (See Figure 9 in the appendix). The remaining two parts are the same as the construction in Figure 2. The following theorem gives a sharper bound on the approximation error compared with Theorem 3.5.
@@
Under the 3.4, if in addition, $\mathbb{P}$ constrained to each component satisfies 3.6, there exists a neural network $\widehat{g}=\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ{\widehat{g}_{1}}$ with the above architecture such that
$W(\widehat{g}_{\#}\lambda,\mathbb{P})\leqslant O\left(L_{1}^{-2/\max_{i}\{d_{i}^{C}\}}+L_{2}^{-2/\max_{i}\{d_{i}^{C}\}}+(\tau-\tau\log\tau)\right),$
where $d_{i}^{C}$ are the dimensions of the connected components in 3.4, $\widehat{g}_{i},i=1,2$ has the form $\left(\widehat{g}_{i}^{1}(x_{1}),\cdots,\widehat{g}_{i}^{N_{C}}(x_{N_{C}})\right)$ and $\widehat{g}_{1}^{j}\in\mathcal{NN}_{1,d_{j}^{C}}(\Theta(d_{j}^{C}),L_{1})$. $\widehat{g}_{2}^{j},\widehat{g}_{3}^{\tau},\tau$ are the same as in Theorem 3.5.
Let $N$ be the number of nonzero weights in a neural network, Theorem 3.8 shows that a deep NN can achieve $O(N^{-2/d})$ error while by Theorem 3.5 a wide network can only achieve $O(N^{-1/d})$ error. Now, we can put things together to construct NCM approximations. For simplicity, we omit the input and output dimensions of the neural network. As we mention in previous sections, our construction (3.3) consists of two parts, $\widehat{f}_{i}$ approximating the structural functions $f_{i}$ in (3.1), and $\widehat{g}_{j}(Z_{j})=\widehat{g}_{3,j}^{\tau}\circ\widehat{g}_{2,j}\circ\widehat{g}_{1,j}(Z_{j})$ approximating the latent variables $U_{j}$. Let $\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0},\mathcal{F}_{1},\mathcal{F}_{2},\tau)$ be a collection of NCMs with structural equation (3.3) and
$\widehat{f}_{i}\in\mathcal{F}_{0},\widehat{g}_{1,j}=(\widehat{g}^{1}_{1,j},\cdots,\widehat{g}^{N_{C,j}}_{1,j}),\widehat{g}^{i}_{1,j}\in\mathcal{F}_{1},\widehat{g}_{2,j}=(\widehat{g}^{1}_{2,j},\cdots,\widehat{g}^{N_{C,j}}_{2,j}),\widehat{g}^{i}_{2,j}\in\mathcal{F}_{2},$
where $N_{C,j}$ is the number of connected components for $U_{j}$ and $\mathcal{F}_{i}$ are function classes.
@@
Given a causal model $\mathcal{M}^{*}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\bm{U})$, suppose that Assumptions 2.3-3.2 hold and the distributions of $U_{C}$ for all $C^{2}$-component satisfy the assumptions of Theorem 3.8. Let $d_{\max}^{\text{in}}$ and $d_{\max}^{\text{out}}$ be the largest input and output dimension of $f_{i}$ in (3.1) and $d_{\max}^{U}$ be the largest dimension of all latent variables. There exists a neural causal model
^$\widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{NN}(\Theta(d^{\text{in}}_{\max}d^{\text{out}}_{\max}),L_{0}),\mathcal{NN}\left(\Theta\left(d^{U}_{\max}\right),L_{1}\right),\mathcal{NN}\left(\Theta\left((d^{U}_{\max})^{2}\right),L_{2}\right),\tau)$
@@
Similar approximation results also hold for wide NN approximation, as presented in Section 3.1. The proof can be easily adapted to the wide NNs architecture.
@@ 4 Consistency of Neural Causal Partial Identification
$F_{n}=\mathop{\mathrm{argmin}}_{\widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n},\mathcal{F}_{2,n},\tau_{n})}$ $\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\widehat{\mathcal{M}}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))],$ (4.1)
@@
where $P^{\mathcal{M}^{*}}_{n},P_{m_{n}}^{\widehat{\mathcal{M}}}$ are the empirical distribution of $P^{\mathcal{M}^{*}},P^{\widehat{\mathcal{M}}}$ with sample size $n,m_{n}$, $\mu_{T}$ is some given measure and $\mathcal{F}_{i,n}$ will be specified later. For example, the counterfactual outcome $\mathbb{E}[Y(1)]$ is a special case of the objective. Our results can be easily generalized to any linear combination of objective functions of this form. We use the Sinkhorn distance because it can be computed efficiently in practice [15]. We want to study if $F_{n}$ gives a useful lower bound as $n\rightarrow\infty$.
To match the observational distribution, we need to increase the width or depth of the NNs we use. As the sample size increases, the number of parameters also increases to infinity, which creates difficulty in the analysis. To obtain consistency, we need to regularize the functions while preserving their approximation power. Surprisingly, if we do not use any regularization, the following proposition implies that consistency may not hold even if the SCM is identifiable.
@@
There exists a constant $c>0$ and an identifiable SCM $\mathcal{M}^{*}$ satisfying Assumptions 2.3-3.6 such that for any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ satisfying $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\epsilon}}(\bm{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
Here, $\mathcal{M}^{*}$ is the ground-truth model and $\mathcal{M}_{\epsilon}$ are the models we use to approximate $\mathcal{M}^{*}$. Proposition 4.1 implies that we need some regularization on $\mathcal{M}_{\epsilon}$. Otherwise, even if the observation distributions are close, their ATEs can be far away. In particular, we may want to regularize the Lipschitz constant of the NN. Much work has been done to impose Lipschitz regularization during the training process [13, 50, 22, 42, 10]. We denote $\text{Lip}(f)$ to be the Lipschitz constant of a function $f$ and define the truncated Lipschitz NN class,
@@
For simplicity, we omit the dimensions and use shorthand $\mathcal{N}\mathcal{N}^{L_{f},K}(W,L)$. The next theorem gives the consistency result of the min estimator. To state the theorem, we define $F_{*}=\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\mathcal{M}^{*}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))]$ to be the true value and $\underline{F}^{L}$ to be the optimal value of the following optimization problem over SCMs with Lipschitz constant $L$.
@@
$\mathcal{F}_{0,n}=$ $\mathcal{NN}^{\widehat{L}_{f},K}\left(W_{0,n},\Theta\left(\log d^{\text{in}}_{\max}\right)\right),\mathcal{F}_{1,n}=\mathcal{NN}^{\infty,\infty}(\Theta(d^{U}_{\max}),L_{i,n}),$
$\mathcal{F}_{2,n}=$ $\mathcal{NN}^{\infty,K}(\Theta((d^{U}_{\max})^{2}),L_{i,n}),$
@@
$s_{n}=O(m_{n}^{-1/\left(d_{\max}^{U}+2\right)}\log m_{n}+\delta_{n}+\log(nm_{n})\lambda_{n}),\quad\delta_{n}=O(n^{-1/{\max\{2,d_{\max}^{U}\}}}\log^{2}(n)).$

### v3

@@ 3 Approximation Error of Neural Causal Models
In this section, we study the expressive power of NCMs, which serves as a key ingredient in proving the consistency result. In particular, given an SCM $\mathcal{M}^{*}$, we want to construct an NCM $\widehat{\mathcal{M}}$ such that the two causal models produce similar interventional results. Unlike in the discrete case balkeCounterfactualProbabilitiesComputational1994 ; xiaCausalNeuralConnectionExpressiveness2022 , latent distributions can be extremely complicated in general cases. The main challenge is how to design the structure of NCMs to ensure strong approximation power.
@@ A SCM $\mathcal{M}$ with causal graph $\mathcal{G}$ has canonical form if
$\widehat{V}_{i}=\begin{cases}\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right),&V_{i}\text{ is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}+\log\left(\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right)\right)_{k}\right\},&V_{i}\text{ is categorical}.\end{cases}$ (3.3)
For the first part, we need to approximate Lipschitz continuous functions. For simplicity, we assume that the domain of the functions are uniform cubes. Similar arguments hold for any bounded cubes. We denote $\mathcal{NN}_{k_{1},k_{2}}(W,L)$ to be the set of ReLu NNs with input dimension $k_{1}$, output dimension $k_{2}$, width $W$ and depth $L$. It has been shown that $\epsilon(\mathcal{NN}_{k_{1},1}(2d_{1}+10,L_{0}),\mathcal{F}_{L}([0,1]^{d_{1}},\mathbb{R}))\leqslant O(L_{0}^{-{2}/{k_{1}}})$ Yarotsky2018OptimalAO , where $\epsilon(\cdot,\cdot)$ denotes the approximation error defined in Section 2. For a vector valued function, we can use a wider NN to approximate each coordinate and get a similar rate.
@@
3.1 Approximating Mixed Distributions by Neural Networks
@@
Figure 1: Architecture of neural network for $4-$dimensional output. The first (yellow) part approximates the distribution on different connected components of the support using the results from perekrestenkoHighDimensionalDistributionGeneration2022a . The width and depth of each block in this part are $W_{1}$ and $L_{1}$. The second (blue) part transforms the distributions on the unit cube to the distributions on the support. The width and depth of each block in the blue part are $W_{2}$ and $L_{2}$. The third (green) part is the Gumbel-Softmax layer. It combines the distributions on different connected components of the support together and outputs the final distribution.
@@
In this subsection, we will extend the results in perekrestenkoHighDimensionalDistributionGeneration2022a to construct a neural network as the push-forward map. It turns out that to get a good approximation of the targeted distribution, the shape of the support is essential. We introduce the following two assumptions for our latent distribution.
@@
3.4 encompasses almost all natural distributions. For example, distributions supported on $[0,1]^{d}$ and closed balls, finitely supported distributions and mixtures of them all satisfy this assumption. 3.4 allows us to transform the support of the targeted distribution into unit cubes and the nice geometric properties of unit cubes facilitate our construction. The next assumption is slightly than 3.4, but we will show later that it can give us a better rate.
@@
3.5 implies that $d\lambda/d(f^{-1}_{\#}\mathbb{P})$ exists and is lowered bounded by a constant $C_{f}$. We show in the appendix (Proposition B.2) that under 3.5, it is possible to simulate any distribution on the unit cubes with Hölder continuous curves and uniform distribution on $[0,1]$. Proposition B.2 can be viewed as an extension of the Skorohod representation theorem.
Now, we briefly explain the construction of the push-forward maps. An example is provided in Section B.1 to illustrate the construction. The NN architecture consists of three parts (see Figure 2). The input dimension is the same as the number of connected components of the support $N_{C}$. For each component $C_{i}$, let $H_{i}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$, where $d_{i}^{C}$ is the dimension of component $C_{i}$ in 3.4. By $\mathbb{P}=(H_{i})_{\#}(H_{i}^{-1})_{\#}\mathbb{P}$ on $C_{i}$, we can approximate $(H_{i}^{-1})_{\#}\mathbb{P}$ first, which is supported on a unit cube. perekrestenkoHighDimensionalDistributionGeneration2022a constructs a wide NN $\widehat{g}$ of width $W$ and constant depth such that $W(\widehat{g}_{\#}\lambda,(H_{i})^{-1}_{\#}\mathbb{P})\leqslant O(W^{-1/d_{i}^{C}})$ where $\lambda$ is the uniform measure on $[0,1]$. Then, we approximate the Lipschitz map $H_{i}$ to $C_{i}$ to pull the distribution back to $C_{i}$. These are the first two parts (yellow and blue blocks in Figure 2) of the architecture.
Suppose that the output of $i$-th coordinate in the first two parts is $\vec{v}_{i}$, the Gumbel-softmax layer in the third part (green box) combines different components of the support. In particular, we want to output $\vec{v}_{i}$ with probability $p_{i}=\mathbb{P}(C_{i})$. Let $V=[\vec{v}_{1},\cdots,\vec{v}_{i}]$, this can be achieved by outputting $VX$, where $X=(X_{1},\cdots,X_{N_{C}})^{\mathsf{T}}$ is a one-hot random vector with $\mathbb{P}({X}_{i}=1)=p_{i}$. To use backpropagation in training, we use the Gumbel-Softmax trick Jang2016CategoricalRW to (approximately) simulate such a random vector, $\widehat{X}^{\tau}_{i}=\frac{\exp((\log p_{i}+G_{i})/\tau)}{\sum_{k=1}^{N_{C}}\exp((\log p_{k}+G_{k})/\tau)},$ where $\tau>0$ is the temperature parameter (a hyperparameter) and $G_{i}\sim\text{Gumbel}(0,1)$ are i.i.d. standard Gumbel variables. As $\tau\rightarrow 0$, the distribution of $\widehat{X}^{\tau}$ converges almost surely to the categorical distribution (maddisonConcreteDistributionContinuous2017, , Proposition 1). In particular, when $\tau=0$, we denote $\widehat{X}_{i}^{0}=\mathbb{I}_{i=\arg\max_{j}\{\log p_{i}+G_{i}\}}$. The output of the last layer is $V\widehat{X}^{\tau}$. Note that the Gumbel-softmax function is differentiable with respect to parameter $\{\log(p_{i})\}_{i=1,\cdots,N_{C}}$. Therefore, we can train the network with common gradient-based algorithms.
In summary, our construction has the form $\widehat{g}=\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}$, where $\widehat{g}_{i},i=1,2$ have the separable form $\left(\widehat{g}_{i}^{1}(x_{1}),\cdots,\widehat{g}_{i}^{N_{C}}(x_{N_{C}})\right)$ and $\widehat{g}_{1}^{j}\in\mathcal{NN}_{1,d_{j}^{C}}(W_{1},L_{1})$, $\ \widehat{g}_{2}^{j}\in\mathcal{NN}_{d_{j}^{C},d}(W_{2},L_{2}),\ j\in[N_{C}]$; $\{d_{j}^{C}\}$ are the dimension of cubes in 3.4. $\widehat{g}_{3}^{\tau}$ is the Gumbel-Softmax layer with temperature parameter $\tau>0$. We slightly abuse the notations and denote this class of neural nets by $\mathcal{NN}(W_{1},L_{1},W_{2},L_{2},\tau)$. Putting things together, we obtain the following theorem.
@@
Given any probability measure $\mathbb{P}$ on $\mathbb{R}^{d}$ that satisfies 3.4, let $\lambda$ be the Lebesgue measure on $[0,1]^{N_{C}}$, where $N_{C}$ is defined in 3.4.
  1. 1.
    There exists a neural network $\widehat{g}\in\mathcal{NN}(W_{1},\Theta(\max_{j}d_{j}^{C}),\Theta(d\cdot\max_{j}d_{j}^{C}),L_{2},\tau)$ with the above architecture (Figure 2) such that
  2. 2.
    Furthermore, if in addition, $\mathbb{P}$ constrained to each component satisfies 3.5, there exists a neural network $\widehat{g}\in\mathcal{NN}(\Theta(\max_{j}d_{j}^{C}),L_{1},\Theta(d\cdot\max_{j}d_{j}^{C}),L_{2},\tau)$ with the above architecture such that
    $W(\widehat{g}_{\#}\lambda,\mathbb{P})\leqslant O\left(L_{1}^{-2/\max_{i}\{d_{i}^{C}\}}+L_{2}^{-2/\max_{i}\{d_{i}^{C}\}}+(\tau-\tau\log\tau)\right).$
The second conclusion of Theorem 3.6 shows that under one additional assumption on the distribution, deep ReLU networks have a stronger approximation power in approximating distributions, which means we can use fewer computational units to achieve the same worst-case theoretical approximation error.
3.2 NCM Construction
Now, we can use the previous distribution approximation results to construct NCM approximations. For simplicity, we omit the input and output dimensions of the neural network. As we mention in previous sections, our construction (3.3) consists of two parts, $\widehat{f}_{i}$ approximating the structural functions $f_{i}$ in (3.1), and $\widehat{g}_{j}(Z_{j})$ approximating the latent variables $U_{j}$. Let $\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0},\mathcal{F}_{1})$ be a collection of NCMs with structural equation (3.3) and $\widehat{f}_{i}\in\mathcal{F}_{0},\widehat{g}\in\mathcal{F}_{1}.$ By combining Theorem 3.3 and Theorem 3.6, we have the following corollary.
@@ 3.2 NCM Construction
Given a causal model $\mathcal{M}^{*}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\bm{U})$, suppose that Assumptions 2.3-3.2 hold and the distributions of $U_{C}$ for all $C^{2}$-component satisfy the assumptions of the second conclusion of Theorem 3.6. Let $d_{\max}^{\text{in}}$ and $d_{\max}^{\text{out}}$ be the largest input and output dimension of $f_{i}$ in (3.1) and $d_{\max}^{U}$ be the largest dimension of all latent variables. There exists a neural causal model
^$\widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{NN}(\Theta(d^{\text{in}}_{\max}d^{\text{out}}_{\max}),L_{0}),\mathcal{NN}(\Theta(d_{\max}^{U}),L_{1},\Theta((d_{\max}^{U})^{2}),L_{2}))$
@@
Similar approximation results also hold for wide NN approximation (the first construction in Theorem 3.6), as presented in Section 3.1. The proof can be easily adapted to the wide NNs architecture. Corollary 4.4 gives an explicit construction of NCM for a given SCM, which shows the expressivity of NCMs for SCM with mixed variables. This result can be viewed as a generalization of the expressivity result in xia2022neural , but our construction is completely different from xia2022neural .
@@ 4 Consistency of Neural Causal Partial Identification
$F_{n}=\mathop{\mathrm{argmin}}_{\widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n})}$ $\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\widehat{\mathcal{M}}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))],$ (4.1)
@@
where $P^{\mathcal{M}^{*}}_{n},P_{m_{n}}^{\widehat{\mathcal{M}}}$ are the empirical distribution of $P^{\mathcal{M}^{*}},P^{\widehat{\mathcal{M}}}$ with sample size $n,m_{n}$, $\mu_{T}$ is a given measure, $S_{\lambda}(\cdot,\cdot)$ is the Sinkhorn distance and $\mathcal{F}_{i,n}$ will be specified later. For example, the counterfactual outcome $\mathbb{E}[Y(1)]$ is a special case of the objective. Our results can be easily generalized to any linear combination of objective functions of this form. We use the Sinkhorn distance because it can be computed efficiently in practice feydy2020geometric . We want to study if $F_{n}$ is a good lower bound as $n\rightarrow\infty$.
To match the observational distribution, we need to increase the width or depth of the NNs we use. As the sample size increases, the number of parameters also increases to infinity, which creates difficulty for our analysis. To obtain consistency, we need to regularize the functions while preserving their approximation power. Surprisingly, if we do not use any regularization, the following proposition implies that consistency may not hold even if the SCM is identifiable.
@@
There exists a constant $c>0$ and an identifiable SCM $\mathcal{M}^{*}$ satisfying Assumptions 2.3-3.5 such that for any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ satisfying $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\epsilon}}(\bm{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
Here, $\mathcal{M}^{*}$ is the ground-truth model and $\mathcal{M}_{\epsilon}$ are the models we use to approximate $\mathcal{M}^{*}$. Proposition 4.1 implies that even if the observation distributions are close and the ground-truth model is identifiable, their ATEs can be far away, regardless of the sample size, which indicates that the method may not be consistent. We need some regularization on $\mathcal{M}_{\epsilon}$ to recover consistency. In particular, we find that regularizing the Lipschitz constant of the NN can help. Much work has been done to impose Lipschitz regularization during the training process cisse2017parseval ; virmauxLipschitzRegularityDeep2018a ; goukRegularisationNeuralNetworks2021 ; pauliTrainingRobustNeural2021 ; bungertCLIPCheapLipschitz2021 . We denote $\text{Lip}(f)$ to be the Lipschitz constant of a function $f$ and define the truncated Lipschitz NN class,
@@
For simplicity, we omit the dimensions and use shorthand $\mathcal{N}\mathcal{N}^{L_{f},K}(W,L)$. Similarly, we also define the class $\mathcal{NN}^{\infty,K}(W_{1},L_{1},W_{2},L_{2},\tau)$ as a truncated version of $\mathcal{NN}(W_{1},L_{1},W_{2},L_{2},\tau)$. The next theorem gives the consistency result of the min estimator. To state the theorem, we define $F_{*}=\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\mathcal{M}^{*}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))]$ to be the true value and $\underline{F}^{L}$ to be the optimal value of the following optimization problem over SCMs with Lipschitz constant $L$.
@@
$\mathcal{F}_{0,n}$ $=\mathcal{NN}^{\widehat{L}_{f},K}\left(W_{0,n},\Theta\left(\log d^{\text{in}}_{\max}\right)\right),\mathcal{F}_{1,n}=\mathcal{NN}^{\infty,K}(\Theta(d^{U}_{\max}),L_{1,n},\Theta((d^{U}_{\max})^{2}),L_{2,n},\tau_{n}),$
@@
$s_{n}=O(m_{n}^{-1/\left(d_{\max}^{U}+2\right)}\log m_{n}+n^{-1/{\max\{2,d_{\max}^{U}\}}}\log^{2}(n)+\log(nm_{n})\lambda_{n}).$